15.03 COST OF TRADE CREDIT AND BANK LOAN

Lancaster Lumber buys $8 million of materials (net of discounts) on terms of 3/5, net 55; and it currently pays on the 5th day and takes discounts. Lancaster plans to expand, which will require additional financing. Assume 365 days in year for your calculations.

  1. If Lancaster decides to forgo discounts, how much additional credit could it obtain? Write out your answer completely. For example, 5 million should be entered as 5,000,000. Round your answer to the nearest cent.
    $________
  2. What would be the nominal cost of that credit? Do not round intermediate calculations. Round your answer to two decimal places.
    ________%
  3. What would be the effective cost of that credit? Do not round intermediate calculations. Round your answer to two decimal places.
    ________ %
  4. If the company could get the funds from a bank at a rate of 8%, interest paid monthly, based on a 365-day year, what would be the effective cost of the bank loan? Do not round intermediate calculations. Round your answer to two decimal places.
    ________ %
  5. Should Lancaster use bank debt or additional trade credit?
    = Bank debt OR Additional trade credit
Answer #1

Given,

Materials = $ 8 million or $ 8000000

Discount rate = 3% or 0.03

Full payment days = 55 days

Discount days = 5 days

Solution :-

1) Additional credit materials (Full Payment - Discount) 365 days days days $8000000 x (55-5) 365 $8000000 365 & so - $109589Effective cost of trade credit 365 so Discount rate (1- Discount rate -7.3 0.03 -0.03 7.3 - + 0.03092783505) - (1.03092783505+ + U.00 0.08 go 12 B 1 + 0.00 6666666667 - (1.00666666666) - = 1.0830 - 1 = 0.0830 of 8.30% Since, effective cost of the ban